# QOSST: A Highly-Modular Open Source Platform for Experimental Continuous-Variable Quantum Key Distribution

### v3

@@ 3.5 Bob
The beat frequency is the frequency difference between the two lasers, and can typically move by a few $\mathrm{MHz}\text{/}$ if the lasers are not very stable. This was not the case in our implementation thanks to the stability of our lasers. The sequence is unshifted by $f_{\mathrm{beat}}$. This step is known as carrier frequency estimation. Then, the Zadoff-Chu sequence is found by cross-correlations, which gives a precise value for the beginning of the sequence. Knowing the number of symbols and the symbol rate, the sequence is reduced to the useful sequence containing only the symbols. The rest of the DSP is performed by subframes because the change in $f_{\mathrm{beat}}$ can be relatively quick and we want to do the analysis on a time scale where $f_{\mathrm{beat}}$ can be considered constant. In particular, this subframe analysis is critical when the lasers used are not very stable. For each subframe, the frequencies of the two pilots are estimated again, and $f_{\mathrm{beat}}$ is also estimated again using Eq. (10). The first pilot is filtered and used as a phase reference, and a matching RRC filter is applied to the quantum data, after it has been shifted to baseband by $f_{\mathrm{beat}}+f_{\mathrm{shift}}$. Following this, the best sampling point is found by the maximal variance (eye diagram) method [34]. The phase is then corrected using the phase reference from the pilot (a uniform filter is applied to the phase reference for better performance). At this point, Alice’s and Bob’s symbols only differ by a global phase (and some noise) and the global phase correction is performed by testing all global rotations and keeping the one with the best covariance term $\langle XY\rangle$. More details on the DSP of both Alice and Bob can be found in appendix E.
@@ 4 Results
In Fig. 6, we plot the excess noise for the experiments using the VOA with attenuation corresponding to $25\;$\mathrm{km}\text{/}$$ and the fiber spool. The estimation of the excess noise is rather stable. The stability of the estimation of $\langle n\rangle$, $T$, shot noise and $V_{el}$ were also checked. The results in terms of the secret key rate are given in Fig. 7.
@@ 5 Discussion
Unfortunately the experiments at $25\;$\mathrm{km}\text{/}$$ do not not give any key rate in the finite-size case with the worst-case estimator method. As was briefly mentioned above, another method is to follow the analysis in [20] for each frame. For $N=10^{6}$, which was the number of symbols in our frame, we unfortunately also got a key rate of zero for each frame. However, increasing $N$ for the analysis, we start to get some frames with a positive finite-size key rate. For the experiment with 25 km, 40 frames out of 100 at $N=10^{7}$ and then all the frames at $N=10^{8},10^{9},10^{10}$. For the experiment with the fiber, there were respectively 14, 77, 93 and 94 frames with a positive finite-size key rate at $N=10^{7},10^{8},10^{9},10^{10}$. We now give the average key rate over all the frames (and in parentheses the average over the positive frames) for the VOA experiment at 25 km: $96\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($240\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $817\;$\mathrm{kbit}\text{/}\mathrm{s}$$, $1.1\;$\mathrm{Mbit}\text{/}\mathrm{s}$$, $1.2\;$\mathrm{Mbit}\text{/}\mathrm{s}$$ for $N=10^{7},10^{8},10^{9},10^{10}$ and for the experiment with the fiber: $39\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($280\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $475\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($617\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $718\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($772\;$\mathrm{kbit}\text{/}\mathrm{s}$)$, $803\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($854\;$\mathrm{kbit}\text{/}\mathrm{s}$$) with the same $N$. This makes us confident that one way to get better results would be to increase the number of points that are used for the estimation of $T$ and $\xi$. We indeed checked that the variance of those two variables has a dependence over the number of samples used to estimate them. This would reduce the confidence interval and give better worst-case estimators and subsequently better key rates.
@@
Since the experiments presented in this paper were done, automatic polarisation compensation with a motorised polarisation controller was implemented and tested, and is released with QOSST.

### v4

@@ 3.5 Bob
The beat frequency is the frequency difference between the two lasers, and can typically move by a few $\mathrm{MHz}\text{/}$ if the lasers are not very stable. This was not the case in our implementation thanks to the stability of our lasers. The sequence is unshifted by $f_{\mathrm{beat}}$. This step is known as carrier frequency estimation. Then, the Zadoff-Chu sequence is found by cross-correlations, which gives a precise value for the beginning of the sequence. Knowing the number of symbols and the symbol rate, the sequence is reduced to the useful sequence containing only the symbols. The rest of the DSP is performed by subframes because the change in $f_{\mathrm{beat}}$ can be relatively quick and we want to do the analysis on a time scale where $f_{\mathrm{beat}}$ can be considered constant. In particular, this subframe analysis is critical when the lasers used are not very stable. In particular the duration of the subframe must be chosen such that the relative frequency variations between the two lasers (including linewidth and change in central frequencies) are less than the invert of the subframe’s duration, during the said subframe. In practice, this can be done by an optimization over the possible subframe sizes as shown in the middle graph of Fig. 5. For each subframe, the frequencies of the two pilots are estimated again, and $f_{\mathrm{beat}}$ is also estimated again using Eq. (10). The first pilot is filtered and used as a phase reference, and a matching RRC filter is applied to the quantum data, after it has been shifted to baseband by $f_{\mathrm{beat}}+f_{\mathrm{shift}}$. Following this, the best sampling point is found by the maximal variance (eye diagram) method [34]. The phase is then corrected using the phase reference from the pilot (a uniform filter is applied to the phase reference for better performance). At this point, Alice’s and Bob’s symbols only differ by a global phase (and some noise) and the global phase correction is performed by testing all global rotations and keeping the one with the best covariance term $\langle XY\rangle$. More details on the DSP of both Alice and Bob can be found in appendix E.
@@ 3.6 Secret Key Rate
Technically, QOSST can emit and receive discrete-modulated sequences (although the DSP does not use in its present form any potential advantage provided by the discrete nature of the constellation). However, QOSST does not yet include methods to compute the SKR in the discrete-modulation scenario. We have planned to add, in future releases of this package, methods following the analysis of Denys and Leverrier [14] for the asymptotic case. The finite-size case is more delicate as, while much progress has been made in the past few years, it is still unclear which analysis will provide the best key rates in practice.
@@ 3.7 Simulations and Post-Processing
3.8 Re-usability of the software
In the hope that our open source platform can be of use to the community, we have undertaken several steps to enhance the re-usability of our code. First, the Hardware Abstraction Layer allows to use the code by only providing the new control commands, and changing a parameter in the configuration file, hence allowing for the use of other pieces of equipment with minimal changes. Second, the behavior of the software is highly configurable with more than 100 parameters in the configuration file, allowing for the adaptation to a variety of situations and optimizations. Finally, the code has been broken down into relatively small functions, allowing their use outside of the main program. For instance, it is easy to use the functions for the DSP, for parameter estimation or for key rate computations, without having to run the full program or use any hardware.
Our software has, however, some limitations in its present form: it cannot be directly used to experiment with time-multiplexed pilots, phase-diverse detection or polarization-diverse setups. While we plan to provide updates to increase the scope of our software in future releases, we hope that the provided documentation is already sufficient for other groups to adapt our software for their experiment, if needed.
@@ 4 Results
In Fig. 6, we plot the excess noise for the experiments using the VOA with attenuation corresponding to $25\;$\mathrm{km}\text{/}$$ and the fiber spool. The estimation of the excess noise is rather stable. The stability of the estimation of $\langle n\rangle$, $T$, shot noise and $V_{el}$ were also checked. The results in terms of the secret key rate are given in Fig. 7. Note that for the fiber spool experiment, only the first 50 frames are shown. Indeed, after this time, the setup drifts away from its optimal position, in particular the polarization of the signal and of the local oscillator start to mismatch, causing a drop in efficiency. This technical issue has now been resolved, as discussed later.
@@ 5 Discussion
Unfortunately the experiments at $25\;$\mathrm{km}\text{/}$$ do not not give any key rate in the finite-size case with the worst-case estimator method. As was briefly mentioned above, another method is to follow the analysis in [20] for each frame. For $N=10^{6}$, which was the number of symbols in our frame, we unfortunately also got a key rate of zero for each frame. However, increasing $N$ for the analysis, we start to get some frames with a positive finite-size key rate. This is done by simply changing $N$ for the correction terms assuming the same estimated values for $\xi$ and $T$, and hence represents what we could get if we were able to obtain the same numbers with larger frames. For the experiment with 25 km, 40 frames out of 100 at $N=10^{7}$ and then all the frames at $N=10^{8},10^{9},10^{10}$. For the experiment with the fiber, there were respectively 14, 77, 93 and 94 frames with a positive finite-size key rate at $N=10^{7},10^{8},10^{9},10^{10}$. We now give the average key rate over all the frames (and in parentheses the average over the positive frames) for the VOA experiment at 25 km: $96\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($240\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $817\;$\mathrm{kbit}\text{/}\mathrm{s}$$, $1.1\;$\mathrm{Mbit}\text{/}\mathrm{s}$$, $1.2\;$\mathrm{Mbit}\text{/}\mathrm{s}$$ for $N=10^{7},10^{8},10^{9},10^{10}$ and for the experiment with the fiber: $39\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($280\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $475\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($617\;$\mathrm{kbit}\text{/}\mathrm{s}$$), $718\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($772\;$\mathrm{kbit}\text{/}\mathrm{s}$)$, $803\;$\mathrm{kbit}\text{/}\mathrm{s}$$ ($854\;$\mathrm{kbit}\text{/}\mathrm{s}$$) with the same $N$. This makes us confident that one way to get better results would be to increase the number of points that are used for the estimation of $T$ and $\xi$. We indeed checked that the variance of those two variables has a dependence over the number of samples used to estimate them. This would reduce the confidence interval and give better worst-case estimators and subsequently better key rates.
@@
Since the experiments presented in this paper were done, automatic polarisation compensation with a motorised polarisation controller was implemented and tested, and is released with QOSST. This automatic compensation algorithm showed stable recovery over tens of hours with a fiber spool.